OpenStudy (anonymous):

In the equation y = 3x − 4, the __________. A. slope is 3, and the y-intercept is 4 B. slope is 3, and the y-intercept is -4 C. slope is -4, and the y-intercept is 3 D. slope is 3, and the x-intercept is -4

OpenStudy (anonymous):

B right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well, this is very important to know! This is a Slope-Intercept Equation y = mx + b Always remember your (m) is your slope, and the (b) is the y-intercept. Slope = m = rise/run = y2 - y1/x2-x1

OpenStudy (anonymous):

Thanks everyone!!